# Adaptive Joint Estimation of Temporal Vertex and Edge Signals

### v2

@@ III-C Joint estimation of vertex and edge signals
AJVEE is targeted at cases where time-varying edge signals influence time-varying vertex signals. We can specify the ALMS-Hodge onto the graph edges by specifying $k=1$, resulting in $\mathbf{H}_{1}=\sum_{p=0}^{P}\hat{\theta}_{1,p}\mathbf{L}_{1}^{p}$. Using Chebyshev approximation, the AJVEE update based on the ALMS-Hodge on edge signals can then be obtained as:
@@
To define the AJVEE expression on the graph vertices, we specify $k=0$ in the ALMS-Hodge, but what is different here on the graph vertices is that, AJVEE will be following our time-varying regression data model defined in Section III-A. It should be highlighted again that the vertex signal $\boldsymbol{x}_{0}[t]$ is under the influence of the time-varying edge signal $\boldsymbol{x}_{1}[t]$. So, the regression parameters are the current step edge signal estimate $\hat{\boldsymbol{x}}_{1}[t]$ obtained at the previous time step $t-1$. As a result, we shall replace the matrix $\mathbf{H}_{k}$ in ALMS-Hodge in (16) with the time-varying vertex regression $\hat{\mathbf{H}}_{0}[t]=\sum_{p=0}^{P}\hat{\theta}_{0,p}(\hat{\mathbf{L}}_{0}[t])^{p}$ to incorporate the influence of the time-varying edge signal $\boldsymbol{x}_{1}[t]$ onto the time-varying vertex signals $\boldsymbol{x}_{0}[t]$. Then, to conduct an estimation on $\boldsymbol{x}_{0}[t]$ when it is under the influence of $\boldsymbol{x}_{1}[t]$, the vertex update strategy in AJVEE based on ALMS-Hodge is
@@
Instead of separately estimating the vertex and edge signals, a joint estimation can be done by the simultaneous operation of (23) and (24). We define a joint shift operator $\mathbf{L}_{J}[t]$ to incorporate the adjacent relationship within one simplex order together with the incidence relationship between different orders of simplicity. Assuming that the signals are on the vertices (0-simplex) and edges (1-simplex), and also assuming that we have the triangle (2-simplex) adjacency but not the 2-simplicial signals, we can formulate such joint shift operator on the error terms $\boldsymbol{e}_{0}[t]$ and $\boldsymbol{e}_{1}[t]$ by
@@
The AJVEE operates in an online fashion where the time-varying edge signal and the time-varying vertex signal are estimated jointly using (28) at each time instance. It should be emphasized that the AJVEE discussed here is not simply estimating edge weights from graph vertex signals nor a link prediction task; the graph edges are regarded as the 1-simplex with signals existing on the edges. The complete AJVEE algorithm is summarized in Algorithm 1.
@@
III-D Spectral analysis of AJVEE
IV Generalizing AJVEE to higher-ordered simplicial complexes
In this section, the AJVEE is defined given a more generalized scenario where the time-varying simplicial signals on $\mathcal{X}_{k}$ are influenced by the time-varying signals simplicial signals on $\mathcal{X}_{k+1}$. Given the topology, we can rely on defining the ALMS-Hodge on $\mathcal{X}_{k}$ and $\mathcal{X}_{k+1}$ shown in Section III-B. First, the $k$-simplex regression in (4) will need to have the parameters defined using the $k+1$-simplex. Then, we can define the joint terms
@@ IV Generalizing AJVEE to higher-ordered simplicial complexes
$\begin{split}&\mathbf{H}_{J}[t]=\text{blkdiag}(\mu_{k}\hat{\mathbf{H}}_{k}[t],\mu_{k+1}\mathbf{H}_{k+1}),\\ &\mathbf{R}_{J,\mathcal{S}}=\text{blkdiag}(r_{k,\mathcal{S}}\mathbf{R}_{k,\mathcal{S}},r_{{k+1},\mathcal{S}}\mathbf{R}_{{k+1},\mathcal{S}})\text{ and }\\ &\mathbf{R}_{J,\bar{\mathcal{S}}}=\text{blkdiag}(r_{k,\bar{\mathcal{S}}}\mathbf{R}_{k,\bar{\mathcal{S}}},r_{{k+1},\bar{\mathcal{S}}}\mathbf{R}_{{k+1},\bar{\mathcal{S}}}),\end{split}$ (32)
@@ V Experiment Results and Discussion
V-A Experiment set up
@@
The datasets we conduct the experiments on are the Sioux Falls transportation network, the Anaheim transportation network, and the England COVID-19 dataset. The edge signal missing rates for the Sioux Falls network and the Anaheim network are $26\%$ and $42\%$ respectively. The England COVID-19 dataset is a dynamic network that changes over time. In the experiment, we perceive every newly appeared edge at the topology in time $t$ compared to the previous topology at $t-1$ as an edge with signals missing, meaning that the corresponding diagonal entry in $\mathbf{D}_{1,\mathcal{S}}$ is 00. All the experiments are conducted in MATLAB 2022a and are repeated $N=100$ times unless specified otherwise. In the experiment, all the tested algorithms will use lowpass filters. The diffusion initialization strategy proposed in [16] is used to initialize the signals for all the experiments. All algorithms running on the same task will have the same filters with the same passband to ensure the fairness of comparison. The performance of all the experiments is measured using the normalized mean squared error (NMSE):
$\text{NMSE}_{k}[t]=\frac{1}{N}\sum_{i=1}^{N}\frac{||(\hat{\boldsymbol{x}}_{k}[t]_{i}-\boldsymbol{x}_{k}[t])||^{2}_{2}}{||(\boldsymbol{x}_{k}[t])||^{2}_{2}},$ (33)
where $||(\hat{\boldsymbol{x}}_{k}[t]_{i}-\boldsymbol{x}_{k}[t])||^{2}_{2}$ is the squared error on the $i^{th}$ run.
The topology for the Sioux Falls transportation network and the Anaheim transportation network are gathered from the real world [52]. The Sioux Falls network consists of $N_{0}=24$ vertices and $N_{1}=38$ edges and the Anaheim network has $N_{0}=406$ vertices and $N_{1}=624$ edges. These two transportation networks have provided static traffic flows on the edges for the traffic under equilibrium [52]. As for the England COVID-19 dataset, it is a series of 61 dynamic graphs [43]. All 61 graphs have $N_{0}=129$ vertices, with a different number of edges $N_{1}[t]$ on each day. The vertices represent different regions in England, recording the number of COVID-19 cases in the region each day, and the edges represent population mobility between regions; self-loops represent the mobility within the region. Figure 7 displays the England COVID-19 dataset at three different time points. Even though there is no data on the $2$-simplex in any of the datasets, the upper adjacency of the edges can still be represented in $\mathbf{L}_{1,u}$ because the topological structures still exist.
V-B ALMS-Hodge Analysis
Since the ALMS-Hodge is a component within AJVEE, we first conduct a few experiments to assess the performance of our ALMS-Hodge procedure in (16) at estimating a time-varying edge signal in an online procedure on the Sioux Falls network. To specify (16) onto the edges, we will be setting $k=1$, resulting in the edge signal estimation formula shown in (23). The ALMS-Hodge is tested under the following two experiment settings. First, fix the noise level and run the ALMS-Hodge for 4 different step sizes. Second, fix the step size and run the ALMS-Hodge for 4 different noise levels. The edge signals are generated synthetically on the Sioux Falls network and the Anaheim network by combining several sinusoidal functions to make the static signal time-varying:
$\boldsymbol{x}_{1}[t]=\sum^{P}_{p=1}\boldsymbol{a}_{p}f_{p}[t],$ (34)
[…]
@@ V-B ALMS-Hodge Analysis
For the fixed noise experiment, the noise is Gaussian distribution with mean $=0$ and variance $=0.1$; four different step sizes are $\mu_{1}=0.1,0.6,1.2$, and $1.8$. The signal reconstruction of the signal on one edge and the NMSE on the entire topology is shown in Figure. 2(a) and Figure. 2(b) respectively. From the edge signal estimation in Figure. 2(a), we can see that the step size parameter $\mu_{1}$ indeed controls the magnitude of the update. For $\mu_{1}=0.1$, the magnitude of the update is too small, which is reflected by visually inspecting Figure. 2(a): the estimation does not catch up with the time-varying change in the signal. As for the case when $\mu_{1}$ is too large ($\mu_{1}=1.8$), the large magnitude causes the estimation to fluctuate and unstable, which can be observed from Figure. 2(a) as well. This confirms that $\mu_{1}$ controls the amount of update of the ALMS-Hodge. Inspecting the NMSE in Figure. 2(b), we can gain some insights on the design choice of $\mu_{1}$. The error of the estimation will be large when $\mu_{1}$ is either too large or too small.
(a) Reconstruction of a missing edge signal.
@@ V-C Joint Estimation on Transportation Network
Returning to the question of joint estimation, we will be testing AJVEE based on Algorithm 1 at jointly estimating time-varying vertex and edge signals by utilizing AJVEE within the ALMS-Hodge on the Sioux Falls network and the Anaheim network. The edge signal generation and missing data masking follow the same procedure in V-B. The ground truth vertex signals are synthetically generated with additive i.i.d. Gaussian noise $\epsilon_{0}[t]$ with zero mean and variance $=0.1$ on both the Sioux Falls network and the Anaheim network:
$\boldsymbol{x}_{0}[t+1]=\hat{\mathbf{H}}_{0}[t]\boldsymbol{x}_{0}[t]+\epsilon_{0}[t],$ (35)
We should point out that even though the vertex signal is synthetic, which seems a bit abstract, it is built using properties extracted from the underlying real-world topology. In the context of a transportation network, one may perceive the vertex signal as a measurement of congestion at an intersection or a number of the crowd at a place of interest. The magnitude of the vertex signal will be proportional to the amount of traffic on the edges. Using AJVEE, we aim to predict the congestion or population on the vertices using the predicted traffic on the edges. Since this paper is not focused on transportation network models, we will not further discuss the meaning of vertex signals and use it purely as the target signal to be estimated. Figure. 1 illustrates an example of such a time-varying vertex signal under the influence of time-varying edge signals on the Sioux Falls network.
@@
To show the advantages of using a TSP approach (the ALMS-Hodge) in AJVEE, we will compare AJVEE against a non-TSP online edge signal estimation named Line Graph LMS (LGLMS) from [53], which uses line graphs to map edge signals on the vertices then processed by GSP. The aggregations terms used by AJVEE in this experiment are (17) for edge signals and (18) for vertex signals. Since the signal aggregation terms are missing in the LGLMS, we added the same signal aggregation terms that we use in AJVEE to LGLMS. For edge signal estimation baselines, in addition to the LGLMS algorithm, we also included a non-adaptive TSP baseline using a basic low pass simplicial filter $\mathbf{\Sigma}_{1,lp}$ similar to (3): $\hat{\boldsymbol{x}}_{1}[t+1]=\mathbf{U}_{1}\mathbf{\Sigma}_{1,lp}\mathbf{U}_{1}^{T}\hat{\boldsymbol{x}}_{1}[t].$ The vertex estimation results of AJVEE will be compared with the following three GSP baselines in addition to the LGLMS: the GLMS algorithm [4], the adaptive Graph Least Mean $p^{th}$ algorithm (GLMP) [12], and the adaptive Graph-Sign (G-Sign) algorithm[11]. Note that by the algorithm definitions in their original literature, the GSP methods are using a fixed graph Laplacian, so among all the baselines only the LGLMS algorithm is aware that the time-varying edge signals influence the vertices. The step sizes and the aggregation weights are tuned using grid search; the final parameter selections for AJVEE are shown in Table I. To promote fair comparison, all the tested algorithms share the same $\mu_{0}$ for vertex signal estimation and all the tested algorithms share the same $\mu_{1}$ for edge signal estimation.
@@
Figure 4: Estimation on one of the edges (top) and one of the vertices (bottom) of the Sioux Falls network.
@@
The NMSE of the signal estimation results over the 100 runs are recorded in Fig. 5 for the Sioux Falls network and in Fig. 6 for the Anaheim network. Additionally, an illustration of the estimation results on one of the missing edges and one of the vertices in the Sioux Falls network is shown in Fig. 4. From Fig. 5 and Fig. 6, AJVEE has lower NMSE than the two baselines for the majority of the time points. From the low NMSE, we can confirm that the AJVEE is accurate at tracking the changes of time-varying edge signals. AJVEE outperforms the LGLMS algorithm because the LGLMS does not operate directly on the edge signal; the LGLMS algorithm treats the original edge signal as vertex signals on a line graph projection. However, the projected signal in LGLMS is not guaranteed to be smooth on the 00-simplices, so using the GSP nodal operations to predict edge signals based on a smooth graph assumption on the line graph projection does not perform well [29]. As for the TSP non-adaptive method, the filtering operation $\mathbf{U}_{1}\mathbf{\Sigma}_{1,lp}\mathbf{U}_{1}^{T}\boldsymbol{y}_{1}[t]$ is not optimized to reduce time-dependent errors compared to the adaptive update $\hat{\mathbf{H}}_{1}\boldsymbol{e}_{1}[t]$, which lead to the less compelling results compared to AJVEE.
TABLE I: Parameters settings
vertex Parameters $\mu_{0}$ $r_{0}$ $\mathcal{F}_{0}$ $P$
Sioux Falls 0.45 0.0005 $[0,0.4\lambda_{0,max}]$ 7
Anaheim 0.75 0.05 $[0,0.4\lambda_{0,max}]$ 7
@@
Sioux Falls 0.45 0.05 0.2 $[0,0.58\lambda_{1,max}]$ 7
Anaheim 0.35 0.001 0.002 $[0,0.2\lambda_{1,max}]$ 7
The AJVEE has the lowest estimation error for the vertex signal estimation shown in Fig. 5 and Fig. 6. The GSP approaches do not quite match up to the performance of AJVEE at predicting the time-varying vertex signals because the GSP methods used a fixed graph Laplacian to process the signals on the vertices. On the other hand, AJVEE utilizes the tive-varying edge signal to represent the diffusion dynamics of the vertex signals, which is more accurate at tracking the temporal evolving patterns of the time-varying vertex signals. Our proposed AJVEE framework in Algorithm 1 has demonstrated the ability to effectively capture the time-varying edge signals jointly with the time-varying vertex signals. LGLMS takes a similar approach as AJVEE to use edge signals to represent the time-varying regression, but the relatively inaccurate edge signal estimation causes the LGLMS to have inaccurate vertex signal estimation.

### v3

@@ IV-B Joint estimation of vertex and edge signals
AJVEE is targeted at cases where time-varying edge signals influence time-varying vertex signals. We can specify the ALMS-Hodge onto the graph edges by specifying $k=1$, resulting in $\mathbf{H}_{1}=\sum_{p=0}^{P}\hat{\theta}_{1,p}\mathbf{L}_{1}^{p}$. Using Chebyshev approximation 7, the AJVEE update based on the ALMS-Hodge on the edge signals can then be obtained as:
@@
To define the AJVEE expression on the graph vertices, we specify $k=0$ in the ALMS-Hodge. However, AJVEE will be following our time-varying regression data model defined in Section IV-A on the graph vertices. The vertex signal $\boldsymbol{x}_{0}[t]$ is under the influence of the time-varying edge signal $\boldsymbol{x}_{1}[t]$ so the regression parameters are the current step edge signal estimate $\hat{\boldsymbol{x}}_{1}[t]$ obtained at the previous time step $t-1$. As a result, the matrix $\mathbf{H}_{k}$ in ALMS-Hodge in (14) is replaced with the time-varying vertex regression $\hat{\mathbf{H}}_{0}[t]=\sum_{p=0}^{P}\hat{\theta}_{0,p}(\hat{\mathbf{L}}_{0}[t])^{p}$ to incorporate the influence of the time-varying edge signal $\boldsymbol{x}_{1}[t]$ onto the time-varying vertex signals $\boldsymbol{x}_{0}[t]$. The vertex update strategy in AJVEE based on ALMS-Hodge is
@@
Instead of separately estimating the vertex and edge signals, a joint estimation can be done by the simultaneous operation of (28) and (29). We define a joint operator $\mathbf{H}_{J}[t]$ to incorporate the adjacent relationship within one simplex order together with the incidence relationship between different orders of simplicity. Assuming that the signals are on the vertices (0-simplex) and edges (1-simplex), and also assuming that we have the triangle (2-simplex) adjacency but not the 2-simplicial signals, we can formulate such joint shift operator on the error terms $\boldsymbol{e}_{0}[t]$ and $\boldsymbol{e}_{1}[t]$ by
@@
The AJVEE operates in an online fashion where the time-varying edge signal and the time-varying vertex signal are estimated jointly using (33) at each time instance. AJVEE simultaneously performs missing signal imputation as it makes predictions because of the incorporation of simplicial diffusion. Additionally, from the spectral filter aspect, the simplicial filter within AJVEE ensures denoising is done along the update. These advantages and characteristics of AJVEE enable AJVEE to overcome the challenges of representing topological irregularity, estimating unknown data from known data, removing noise present in the data, and extracting cross-space-time variations we pointed out at the beginning of Section I. The complete AJVEE algorithm is summarized in Algorithm 1. For the spectral analysis of the AJVEE from the TSP perspective, please refer to Appendix B.
@@
IV-C Generalizing AJVEE to high-ordered simplicial complexes
In this section, the AJVEE is defined given a more generalized scenario where the time-varying simplicial signals on $\mathcal{X}_{k}$ are influenced by the time-varying signals simplicial signals on $\mathcal{X}_{k+1}$. Given the topology, we can rely on defining the ALMS-Hodge on $\mathcal{X}_{k}$ and $\mathcal{X}_{k+1}$ shown in Section III. First, the $k$-simplex regression in (22) will need to have the parameters defined using the $k+1$-simplex. Then, we can define the joint terms
@@ IV-C Generalizing AJVEE to high-ordered simplicial complexes
$\begin{split}&\mathbf{H}_{J}[t]=\text{blkdiag}(\mu_{k}\hat{\mathbf{H}}_{k}[t],\mu_{k+1}\mathbf{H}_{k+1}),\\ &\mathbf{R}_{J,\mathcal{S}}=\text{blkdiag}(r_{k,\mathcal{S}}\mathbf{R}_{k,\mathcal{S}},r_{{k+1},\mathcal{S}}\mathbf{R}_{{k+1},\mathcal{S}})\text{ and }\\ &\mathbf{R}_{J,\bar{\mathcal{S}}}=\text{blkdiag}(r_{k,\bar{\mathcal{S}}}\mathbf{R}_{k,\bar{\mathcal{S}}},r_{{k+1},\bar{\mathcal{S}}}\mathbf{R}_{{k+1},\bar{\mathcal{S}}}).\end{split}$ (35)
@@ V Experiment Results and Discussion
V-A Experiment settings
The datasets we conduct the experiments on are the Sioux Falls transportation network, the Anaheim transportation network, and the England COVID-19 dataset. All the experiments are conducted in MATLAB 2022a and are repeated $N=100$ times unless specified otherwise. In the experiment, all the tested algorithms will use lowpass filters defined in (8). The diffusion initialization strategy proposed in [30] is used to initialize the signals for all the experiments. All algorithms running on the same task will have the same filters with the same passband to ensure the fairness of comparison. The performance of all the experiments is measured using the normalized mean squared error (NMSE):
$\text{NMSE}_{k}[t]=\frac{1}{N}\sum_{i=1}^{N}\frac{\|(\hat{\boldsymbol{x}}_{k}[t]_{i}-\boldsymbol{x}_{k}[t])\|^{2}_{2}}{\|(\boldsymbol{x}_{k}[t])\|^{2}_{2}},$ (36)
where $\|(\hat{\boldsymbol{x}}_{k}[t]_{i}-\boldsymbol{x}_{k}[t])\|^{2}_{2}$ is the squared error on the $i^{th}$ run.
The topology for the Sioux Falls transportation network and the Anaheim transportation network are gathered from the real world [44]. The Sioux Falls network consists of $N_{0}=24$ vertices and $N_{1}=38$ edges and the Anaheim network has $N_{0}=406$ vertices and $N_{1}=624$ edges. These two transportation networks have provided static traffic flows on the edges for the traffic under equilibrium [44]. The vertex and edge signal missing rates for the Sioux Falls network is $26\%$ and for the Anaheim network is $30\%$. The missing signal observations in the Sioux Falls network are modeled using the spectral sampling strategy discussed in [6] and the missing signal observations in the Anaheim network are set to be random for each experiment run but will remain fixed within an experiment run. The edge signals are generated synthetically by combining several sinusoidal functions to make the static signal provided by the dataset time-varying:
$\boldsymbol{x}_{1}[t]=\sum^{P}_{p=1}\boldsymbol{a}_{p}f_{p}[t],$ (37)
where the signals $\boldsymbol{a}_{p}$ are static edge signals and the term $f_{p}[t]$ is a sinusoidal function. The ground truth vertex signals are synthetically generated with additive i.i.d. Gaussian noise $\epsilon_{0}[t]$ with zero mean and standard deviation of $=0.2$ on both the Sioux Falls network and the Anaheim network:
$\boldsymbol{x}_{0}[t+1]=\hat{\mathbf{H}}_{0}[t]\boldsymbol{x}_{0}[t]+\epsilon_{0}[t].$ (38)
We should point out that even though the vertex signal is synthetic, which seems a bit abstract, it is built using properties extracted from the underlying real-world topology. In the context of a transportation network, one may perceive the vertex signal as a measurement of congestion at an intersection or a number of the crowd at a place of interest. The magnitude of the vertex signal will be proportional to the amount of traffic on the edges. Using AJVEE, we aim to predict the congestion or population on the vertices using the predicted traffic on the edges. Since this paper is not focused on transportation network models, we will not discuss the meaning of vertex signals further and will use them purely as the target signal to be estimated. Figure. 1 illustrates an example of such a time-varying vertex signal under the influence of time-varying edge signals on the Sioux Falls network.
The England COVID-19 dataset consists of a series of 61 dynamic graphs [36]. In the experiment, we perceive every newly appeared edge at the topology in time $t$ compared to the previous topology at $t-1$ as an edge with signals missing, meaning that the corresponding diagonal entry in $\mathbf{D}_{1,\mathcal{S}}$ is 00. All 61 graphs have $N_{0}=129$ vertices, with a different number of edges $N_{1}[t]$ on each day. The vertices represent different regions in England, recording the number of COVID-19 cases in the region each day, and the edges represent population mobility between regions; self-loops represent the mobility within the region. The upper row of Figure 7 displays the England COVID-19 dataset at three different time points after our preprocessing. Even though there is no data on the $2$-simplex in any of the datasets, the upper adjacency of the edges can still be represented in $\mathbf{L}_{1,u}$ because the topological structures still exist.
@@ V-A Experiment settings
(a) Reconstruction of an unobserved edge signal.
@@
V-B ALMS-Hodge analysis
Since the ALMS-Hodge is a component within AJVEE, we first conduct a few experiments to assess the performance of our ALMS-Hodge procedure in (14) at estimating a time-varying edge signal in an online procedure on the Sioux Falls network. To specify (14) onto the edges, we will be setting $k=1$, resulting in the edge signal estimation formula shown in (28). The ALMS-Hodge is tested under the following two experiment settings. First, fix the noise level and run the ALMS-Hodge for 4 different step sizes. Second, fix the step size and run the ALMS-Hodge for 4 different noise levels.
[…]
For the fixed noise experiment, the noise is Gaussian distribution with mean $=0$ and variance $=0.1$; four different step sizes are $\mu_{1}=0.1,0.6,1.2$, and $1.8$. The signal reconstruction of the signal on one edge and the NMSE on the entire topology is shown in Figure. 2(a) and Figure. 2(b) respectively. The NMSE of comparing the ground truth against the noisy containing missing edge signal observations is included in Figure 2(a) and denoted as noisy&\&&missing. From the edge signal estimation in Figure. 2(a), we can see that the step size parameter $\mu_{1}$ indeed controls the magnitude of the update. For $\mu_{1}=0.1$, the magnitude of the update is too small, which is reflected by visually inspecting Figure. 2(a): the estimation does not catch up with the time-varying change in the signal. As for the case when $\mu_{1}$ is too large ($\mu_{1}=1.8$), the large magnitude causes the estimation to fluctuate and unstable, which can be observed from Figure. 2(a) as well. This confirms that $\mu_{1}$ controls the amount of update of the ALMS-Hodge. The error of the estimation will be large when $\mu_{1}$ is either too large or too small. Inspecting the NMSE in Figure. 2(b), we can gain some insights on the design choice of $\mu_{1}$.
@@ V-C Joint Estimation on transportation network
Returning to the question of joint estimation, we will be testing AJVEE based on Algorithm 1 at jointly estimating time-varying vertex and edge signals by utilizing AJVEE within the ALMS-Hodge on the Sioux Falls network and the Anaheim network.
@@
To show the advantages of using a TSP approach (the ALMS-Hodge) in AJVEE, we created a comparison algorithm that uses Line Graph LMS [45] instead of using the ALMS-Hodge in the AJVEE. We denote this approach as LGLMS in later experiments. Different from the TSP-based ALMS-Hodge, the LGLMS uses line graphs to map edge signals on the vertices then processed by GSP. The aggregations terms used by AJVEE in this experiment are (16) for edge signals and (17) for vertex signals. Since the signal aggregation terms are missing in the LGLMS, we added the same signal aggregation terms that we use in AJVEE to LGLMS. For edge signal estimation baselines, in addition to the LGLMS algorithm, we also included a non-adaptive TSP baseline using a basic low pass simplicial filter $\mathbf{\Sigma}_{1,\mathcal{F}_{lp}}$ similar to (3): $\hat{\boldsymbol{x}}_{1}[t+1]=\mathbf{U}_{1}\mathbf{\Sigma}_{1,\mathcal{F}_{lp}}\mathbf{U}_{1}^{T}\hat{\boldsymbol{x}}_{1}[t]$. The last edge signal baseline is a Moving Average (MA) of the past 5 observations.
The vertex estimation results of AJVEE will be compared with the following three GSP baselines: the GLMS algorithm [4], the adaptive Graph Least Mean $p^{th}$ algorithm (GLMP) [26], and the adaptive Graph-Sign (G-Sign) algorithm[29]. Note that by the algorithm definitions in their original literature, the GSP methods are using a fixed graph Laplacian, so among all the baselines only the LGLMS algorithm is aware that the time-varying edge signals influence the vertices. In addition, we will feed the edge signal estimation of LGLMS into the vertex estimation scheme of the ALMS-Hodge as a comparison to purely using the ALMS-Hodge in AJVEE. The MA is also included for the vertex signal estimation.
The step sizes $\mu_{k}$ and the aggregation weights $r_{k}$ are tuned using grid search. As for the filter used in the experiment, we again use an ideal low-pass filter defined by (8) on both the vertices and the edges. The final parameter selections for AJVEE are shown in Table II. To promote fair comparison, all the tested algorithms share the same $\mu_{0}$ for vertex signal estimation and all the tested algorithms share the same $\mu_{1}$ for edge signal estimation. Similar principle applies to the AJVEE and LGLMS for the aggregation weights. The NMSE of the signal estimation results over the 100 runs are recorded in Fig. 5 for the Sioux Falls network and in Fig. 6 for the Anaheim network. Additionally, an illustration of the estimation results on one of the unobserved edges and one of the unobserved vertices in the Sioux Falls network is shown in Fig. 4.
Refer to captionRefer to caption
@@
Figure 4: Estimation on one of the unobserved edges (top) and one of the unobserved vertices (bottom) of the Sioux Falls network.
@@
From Fig. 5 and Fig. 6, AJVEE has lower NMSE than the baselines for the majority of the time points for the edge estimation task. From the low NMSE, we can confirm that the AJVEE is accurate at tracking the changes of time-varying edge signals. AJVEE outperforms the LGLMS algorithm because the LGLMS does not operate directly on the edge signal; the LGLMS algorithm treats the original edge signal as vertex signals on a line graph projection. However, the projected signal in LGLMS is not guaranteed to be smooth on the 00-simplices, so using the GSP nodal operations to predict edge signals based on a smooth graph assumption and a low-pass filter on the line graph projection does not perform well [46]. As for the TSP non-adaptive method, the filtering operation $\mathbf{U}_{1}\mathbf{\Sigma}_{1,\mathcal{F}_{lp}}\mathbf{U}_{1}^{T}\boldsymbol{y}_{1}[t]$ is not optimized to reduce time-dependent errors compared to the adaptive update $\hat{\mathbf{H}}_{1}\boldsymbol{e}_{1}[t]$, which lead to the less compelling results compared to AJVEE. The MA has a relatively high NMSE because it can not make predictions on the missing edges, which can be observed from Figure 4 that the MA consistently outputs 0 for unobserved signals. The other graph algorithms can infer the missing edge signals because each of the updates implies spatial diffusion, which means that these algorithms conduct spatial missing signal imputation and temporal 1-step predictions at the same time.
TABLE II: Parameters settings
Vertex Parameters $\mu_{0}$ $r_{0,\mathcal{S}}$ $r_{0,\bar{\mathcal{S}}}$ $\mathcal{F}_{0}$ $P$
Sioux Falls 1.25 0.0025 0.05 $[0,0.4\lambda_{0,max}]$ 7
Anaheim 1.1 0.0001 0.00001 $[0,0.4\lambda_{0,max}]$ 7
@@
Sioux Falls 0.45 0.0025 0.15 $[0,0.58\lambda_{1,max}]$ 7
Anaheim 0.75 0.00025 0.0005 $[0,0.58\lambda_{1,max}]$ 7
The AJVEE has the lowest estimation error for the vertex signal estimation shown in Fig. 5 and Fig. 6. The GSP approaches do not quite match the performance of AJVEE in predicting the time-varying vertex signals because the GSP methods used a fixed graph Laplacian to process the signals on the vertices. Here, notice that the LGLMS performs only slightly worse than AJVEE. The reason behind this is because, first, here in the vertex estimation the LGLMS is actually feeding the edge estimation results of LGLMS into the vertex update portion of the AJVEE, so here the LGLMS is essentially another variant of AJVEE. Second, the LGLMS estimation also benefited from the aggregation terms derived for ALMS-Hodge. On the other hand, AJVEE utilizes the tive-varying edge signal to represent the diffusion dynamics of the vertex signals, which is more accurate at tracking the temporal evolving patterns of the time-varying vertex signals. Our proposed AJVEE framework in Algorithm 1 has demonstrated the ability to effectively capture the time-varying edge signals jointly with the time-varying vertex signals. LGLMS takes a similar approach as AJVEE to use edge signals to represent the time-varying regression, but the relatively inaccurate edge signal estimation causes the LGLMS to have inaccurate vertex signal estimation.